# Most Juntas Saturate the Hardcore Lemma

### v1

@@ 1 Introduction
Conceptually, the theorem says that circuit hardness can be explained by a subset of “hard inputs” $H$ on which the function looks random to small circuits. [^1] Holenstein [Hol05] gives a set $H$ of optimal density $2\delta$, but suffers a larger degradation from $s$ to $O\left(\frac{s\gamma^{2}}{n}\right)$. This phenomenon has found applications throughout computer science, including hardness amplification [O’D02, Tre05], pseudorandomness [STV99, CL21], cryptography [Hol05], algorithmic fairness [CDV24], combinatorics [RTTV08], and learning theory [BHK09, KS99].
@@
Such a degradation was shown to be necessary for a certain class of proofs [LTW11], but an unconditional result remained elusive, as proving this theorem appears to require constructing an explicit, mildly approximable function that demands large circuits to strongly approximate it. This felt tantamount to proving breakthrough circuit lower bounds. In a recent work, Blanc, Hayderi, Koch, and Tan [BHKT24] evaded this barrier by arguing about a nonexplicit function. In fact, the proof of Theorem 1 first shows the result for $H\sim\{0,1\}^{n}$ being the weaker notion of a $\delta$-smooth distribution (i.e., no string has more than $\frac{1}{\delta 2^{n}}$ probability of being sampled), and then uses the distribution to extract a density-$\delta$ subset. [BHKT24] proves a tightness result over the more general $\delta$-smooth distributions.
Theorem 2 ([BHKT24, Theorem 2]).
Let $\delta\in(0,1),\gamma\in(0,\tfrac{1}{2})$, $n\in\mathbb{N}$ large enough, and $s\geq n$ with $\Omega\left(\tfrac{1}{\gamma^{2}}\right)\leq s\leq O\left(\tfrac{2^{\gamma^{2}n}}{\gamma^{2}n}\right)$. There exists $f:\{0,1\}^{n}\to\{0,1\}$ such that
@@
The above is a reparametrized version of what appears in [BHKT24], which includes dependencies on $\delta$ from their proof (where $\delta$ was assumed to be constant).222[BHKT24, Theorem 2] gives, for any parameters $s$ and $\gamma$, a tightness result for a function of input length $n(s,\gamma)$. We have reparametrized to the standard convention of fixing the input length to $n$, and examining which $s,\gamma$ are possible (in terms of $n$). Hence, in the regime $\delta=\Theta(1)$ and $s\leq O(2^{\gamma^{2}n}/(\gamma^{2}n))$, [BHKT24] shows that the $(\gamma^{2})$-factor decay in size is tight. Structurally, their argument is very analytically involved and is in multiple stages. In what follows, we say $f$ $\gamma$-correlates with $g$ over $H$ if $\Pr_{x\sim H}[f(x)=g(x)]\geq\frac{1}{2}+\gamma$, and $f$ $\gamma$-approximates $g$ over $H$ if $\Pr_{x\sim H}[f(x)=g(x)]\geq\gamma$.
  • They first prove that for any $\delta$-smooth distribution $H$, the majority on $k$ bits is $\tfrac{\delta}{\sqrt{k}}$-correlated with a $1$-junta over $H$, but no $0.01k$-junta can $\tfrac{1}{4}$-correlate with it over the uniform distribution. [^3] A $k$-junta $f:\{0,1\}^{n}\to\{0,1\}$ is a Boolean function depending only on a subset of $k<n$ input variables. The junta complexity of a function is the smallest $k$ such that the function is a $k$-junta.
  • They then bootstrap this result to show that, for any $\delta$-smooth $H$, the majority of $k$ random functions on $k$ disjoint $\tfrac{n}{k}$-bit input blocks is $\tfrac{\delta}{\sqrt{k}}$-correlated with a size-$O\left(\frac{2^{n/k}}{n/k}\right)$ circuit over $H$, but requires size $\Omega\left(\frac{2^{n}}{n}\right)$ to $\tfrac{1}{4}$-correlate with (over the uniform distribution). They accomplish this by introducing an analytic relaxation of junta complexity, using Fourier-analytic noise-stability arguments to equate this relaxation to the original measure (up to constant factors), and then using coupling arguments, Fourier-analytic calculations, and subgaussian concentration to analyze the relaxed junta complexity of the random-function ensemble.
We note this junta-to-circuit lifting theorem is interesting in its own right, and towards proving this, they establish a novel direct sum theorem. These techniques are also used in [BHKT24] to tightly characterize the sample-complexity overhead of smooth boosting.
@@
Formally, Theorem 2 and Theorem 3 are incomparable: Theorem 2 holds for $s\in[\Omega(1/\gamma^{2}),O(2^{\gamma^{2}n}/(\gamma^{2}n))]$, while Theorem 3 holds for $s\in[\Omega(1/\gamma^{2+\varepsilon}),O(2^{n}/n)]$. While incomparable, we note the former interval is extremely restrictive. For example, the former interval becomes empty in the mild regime of $\gamma=\tfrac{1}{\sqrt{n}}$, but the latter interval remains nearly maximal even when $\gamma=2^{-\Omega(n)}$. The latter interval is short of subsuming the former interval only by an arbitrarily small polynomial factor, $\gamma^{-\varepsilon}$. Removing this $\varepsilon$-slack remains open.
@@
These two points allow our theorem to remain meaningful even for $\delta=o(1)$.
In summary, our Theorem 3 is stronger in the regime $s\geq 1/\gamma^{2+\varepsilon}$, but gives inferior bounds when $1/\gamma^{2}\leq s\leq(1/\gamma)^{2+\varepsilon}$. [^4] For general $s=\Omega(1/\gamma^{2})$, we get approximating circuits of size $O\left(s\gamma^{2}\cdot\frac{\log s}{\log(s\gamma^{2})}\right)$. See Remark 1. We also note that 1 remains open. In particular, it would be interesting to pin down the optimal dependence of the circuit-size decay on $\delta$.
The main technical lemma is a strengthening of a beautiful result of Andreev, Clementi, and Rolim [ACR97], which shows that arbitrary Boolean functions can be approximated by small-size circuits.
@@
This result was proven by Andreev, Clementi, and Rolim [ACR97] in the case where $H$ is uniform over $\{0,1\}^{n}$. A short probabilistic argument proves that the circuit size cannot be improved [ACR97, Theorem 4.1].
@@ 1.1 Proof Overview
Although the weaker version of Theorem 4 suffices, it is our impression that the result of [ACR97] is not as well known to the community as it should be. It is a complete resolution to a very natural question in circuit complexity (it is the “approximate version” of Lupanov’s theorem [Lup71]), and the ideas behind the construction are quite useful (e.g., they appear to have been rediscovered in the construction of covering codes of Rabani and Shpilka [RS10]). This is potentially due to the paper being quite technical and terse, as well as evading search engines. For this reason, we hope to bring attention to this result by giving an exposition and simpler, self-contained proof of Theorem 4 in the appendix, assuming a basic consequence of the fourth-moment method [Ber97] and the existence of asymptotically good codes encodable by linear-size circuits [Spi96]. We now provide a proof overview below.
@@ 1.1.1 Overview of Theorem 4
For ease of exposition, we give an overview of Theorem 4 when $H$ is fixed to be uniform over $\{0,1\}^{n}$. Extending the given arguments to arbitrary distributions $H$ just requires a couple of extra modifications. This overview is morally the same as one provided by Trevisan [Tre09], but perhaps with a slightly different point of view in the latter half.
@@
Unfortunately, a fully random truth table can only be generated by maximally sized circuits. However, we could hope to use a pseudorandom string instead. The only property used about the randomness of $f$ was that its truth table had square-root anticoncentration on subcubes. It turns out $4$-wise uniform strings have square-root anticoncentration with constant probability [Ber97], motivating us to look at this primitive. Implementing the usual 4-wise uniform-generator construction naively in a circuit immediately gives a distribution $\mathcal{C}$ over circuits of size $O(n^{2})$ such that the truth table of $C\sim\mathcal{C}$ is a $4$-wise independent string. With more effort, one can get a distribution over $O(n)$-size circuits, which is optimal. By the fourth-moment method [Ber97], we can argue that the average number of subcubes of $f\oplus\mathcal{C}$ with square-root anticoncentration is at least a constant proportion. Fixing $C\in\mathcal{C}$ that achieves this average, it follows that there is a function $g$ on the first $k$ bits that approximates $f\oplus C$ well by our previous analysis. Consequently, $C\oplus g$ is a circuit of size $O(2^{k}/k+n)$ that approximates $f$ well, as desired.
@@ 2 Preliminaries
Definition 1 ($k$-wise uniformity [HH24]).
@@
$\Pr_{x\sim X}\left[\left|\sum v_{i}x_{i}\right|\geq\frac{\lVert v\rVert_{2}}{\sqrt{3}}\right]\geq\frac{2}{11}.$
@@
This is a $4$-wise uniform generator (see [HH24, Theorem 2.2]). Notice that, as a function of $x$, $G(s)_{x}$ is an evaluation of a degree-$3$ polynomial, which can be done in a circuit of size $O(n^{2})$ by grade-school multiplication (better multiplication algorithms are known, but this suffices). ∎
This theorem suffices to prove Theorem 3. A technical contribution of [ACR97], and a key ingredient behind Theorem 4, is a $4$-wise uniform generator that can be locally computed in linear circuit size, which is the best one could hope for.
Theorem 7 ([ACR97]).
@@ 3 Tightness of Impagliazzo’s Hardcore Lemma
Let $G$ be the $4$-wise uniform generator guaranteed by the hypothesis. By Theorem 5, we have that, for each $c$ and random $s$,
$\Pr_{s\sim\{0,1\}^{m}}\left[\left|\sum_{y\sim\{0,1\}^{\ell}}H_{c}(y)(-1)^{f(c\circ y)+G(s)_{c\circ y}}\right|\geq\sqrt{\frac{\sum H(x)^{2}}{3}}\right]\geq\frac{2}{11}.$
Letting $I_{c}(s)$ denote the indicator of the above event, it follows that
@@
and therefore there exists a choice of $s$ such that $\Pr_{c\sim H^{\prime}}[I_{c}(s)=1]\geq 2/11$. Fix such an $s$. Now define $h:\{0,1\}^{n-\ell}\to\{0,1\}$ by the map
@@
$\geq\frac{2}{11}\cdot\sqrt{\frac{\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)^{2}}{3}}$
$\geq\frac{2}{11}\sqrt{\frac{(1/2^{\ell})\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)}{3}}$
$=\frac{2}{11}\sqrt{\frac{1}{2^{\ell}\cdot 3}}\geq 2\gamma,$
where the penultimate inequality is an application of Cauchy–Schwarz. Hence,
@@
Let $f$ be a random $k$-junta, where $k$ will be chosen soon. In particular, pick a function $g:\{0,1\}^{k}\to\{0,1\}$ uniformly at random, and define $f(x)\coloneqq g(x_{[k]})$ to be $g$ on the first $k$ bits of the input. By a Chernoff bound, we know that for a fixed circuit $C$ and $\delta<0.49$,
$\Pr_{f}\left[\Pr_{x}[f(x)=C(x)]\geq 1-\delta\right]\leq 2^{-\Omega(2^{k})}.$
Set $2^{k}=\Omega(s\log s)$. Taking a union bound over all circuits of size $s$ (of which there are $s^{O(s)}$) we note that, with probability at most $s^{O(s)}2^{-\Omega(2^{k})}<1$, $f$ can be $(1-\delta)$-approximated by a size-$s$ circuit. Fix an $f$ that cannot.
Now consider an arbitrary distribution $H$ over $\{0,1\}^{n}$. The marginal distribution of $H$ on the first $k$ bits will be some distribution $H^{\prime}$ over $\{0,1\}^{k}$. As $\gamma>s^{-1/2}>27\cdot 2^{-k/2}$ for large enough $n$, Lemma 1 and Theorem 6 give a circuit $C:\{0,1\}^{k}\to\{0,1\}$ of size $O\left(\frac{\gamma^{2}2^{k}}{\log(\gamma^{2}2^{k})}+k^{2}\right)$ such that $\Pr_{x\sim H^{\prime}}[g(x)=C(x)]\geq\frac{1}{2}+\gamma.$ Letting $C^{\prime}:\{0,1\}^{n}\to\{0,1\}$ be the circuit that applies $C$ to the first $k$ bits, we note that
$\Pr_{x\sim H^{\prime}}[g(x)=C(x)]=\Pr_{x\sim H}[f(x)=C^{\prime}(x)]\geq\frac{1}{2}+\gamma.$
As $C^{\prime}$ has the same size as $C$, and $\frac{\gamma^{2}2^{k}}{\log(\gamma^{2}2^{k})}+k^{2}=O\left(s\gamma^{2}\cdot\left(\frac{\log s}{\log(\gamma^{2}s)}\right)+\log^{2}s\right)=O_{\varepsilon}(s\gamma^{2})$ when $s\geq 1/\gamma^{2+\varepsilon}$, $C^{\prime}$ is a circuit that $\gamma$-correlates with $f$ over $H$. Hence, $f$ is a $k$-junta that cannot be $(1-\delta)$-approximated by a size-$s$ circuit, but, over any $H$, $\gamma$-correlates with a size-$O_{\varepsilon}(s\gamma^{2})$ circuit, as desired. ∎

### v2

@@ 1 Introduction
Conceptually, the theorem says that circuit hardness can be explained by a subset of “hard inputs” $H$ on which the function looks random to small circuits. [^1] Holenstein [9] gives a set $H$ of optimal density $2\delta$, but suffers a larger degradation from $s$ to $O\left(\frac{s\gamma^{2}}{n}\right)$. This phenomenon has found applications throughout computer science, including hardness amplification [14, 19], pseudorandomness [18, 6], cryptography [9], algorithmic fairness [5], combinatorics [16], and learning theory [2, 11].
@@
Such a degradation was shown to be necessary for a certain class of proofs [12], but an unconditional result remained elusive, as proving this theorem appears to require constructing an explicit, mildly approximable function that demands large circuits to strongly approximate it. This felt tantamount to proving breakthrough circuit lower bounds. In a recent work, Blanc, Hayderi, Koch, and Tan [4] evaded this barrier by arguing about a nonexplicit function. In fact, the proof of Theorem 1 first shows the result for $H\sim\{0,1\}^{n}$ being the weaker notion of a $\delta$-smooth distribution (i.e., no string has more than $\frac{1}{\delta 2^{n}}$ probability of being sampled), and then uses the distribution to extract a density-$\delta$ subset. [4] proves a tightness result over the more general $\delta$-smooth distributions.
Theorem 2 ([4, Theorem 2]).
Let $\delta\in(0,1),\gamma\in\left[\Omega\left(\tfrac{1}{\sqrt{n}}\right),\tfrac{1}{2}\right)$, $n\in\mathbb{N}$ large enough. For $s\in[\Omega(\tfrac{1}{\gamma^{2}}),O(\frac{2^{\gamma^{2}n}}{\gamma^{4}n})]$, there exists $f:\{0,1\}^{n}\to\{0,1\}$ such that
@@
The above is a reparametrized version of what appears in [4], which includes dependencies on $\delta$ from their proof (where $\delta$ was assumed to be constant).222[4, Theorem 2] gives, for any parameters $s$ and $\gamma$, a tightness result for a function of input length $n(s,\gamma)$. We have reparametrized to the standard convention of fixing the input length to $n$, and examining which $s,\gamma$ are possible (in terms of $n$). Hence, in the regime $\delta=\Theta(1)$, $\gamma\geq\Omega(1/\sqrt{n})$ and $s=O(2^{\gamma^{2}n}/(\gamma^{4}n))$, [4] shows that the $\gamma^{2}$-factor decay in size is tight. Structurally, their argument is very analytically involved and is in multiple stages. In what follows, we say $f$ $\gamma$-correlates with $g$ over $H$ if $\Pr_{x\sim H}[f(x)=g(x)]\geq\frac{1}{2}+\gamma$, and $f$ $\gamma$-approximates $g$ over $H$ if $\Pr_{x\sim H}[f(x)=g(x)]\geq\gamma$.
  • Let $k=1/\gamma^{2}$. They first prove that for any $\delta$-smooth distribution $H$, the majority on $k$ bits is $\tfrac{\delta}{\sqrt{k}}$-correlated with a $1$-junta over $H$, but no $0.01k$-junta can $\tfrac{1}{4}$-correlate with it over the uniform distribution. [^3] A $k$-junta $f:\{0,1\}^{n}\to\{0,1\}$ is a Boolean function depending only on a subset of $k<n$ input variables. The junta complexity of a function is the smallest $k$ such that the function is a $k$-junta.
  • They then bootstrap this result to show that, for any $\delta$-smooth $H$, the majority of $k$ random functions on $k$ disjoint $\tfrac{n}{k}$-bit input blocks is $\tfrac{\delta}{\sqrt{k}}$-correlated with a size-$O\left(\frac{2^{n/k}}{n/k}\right)$ circuit over $H$, but requires size $\Omega\left(\frac{k2^{n/k}}{n/k}\right)$ to $\tfrac{1}{4}$-correlate with (over the uniform distribution). They accomplish this by introducing an analytic relaxation of junta complexity, using Fourier-analytic noise-stability arguments to equate this relaxation to the original measure (up to constant factors), and then using coupling arguments, Fourier-analytic calculations, and subgaussian concentration to analyze the relaxed junta complexity of the random-function ensemble.
We note this junta-to-circuit lifting theorem is interesting in its own right, and towards proving this, they establish a novel direct sum theorem. These techniques are also used in [4] to tightly characterize the sample-complexity overhead of smooth boosting.
@@
Formally, Theorem 2 and Theorem 3 are incomparable: Theorem 2 holds only for $\gamma\geq\Omega(1/\sqrt{n})$ and $s\in[\Omega(1/\gamma^{2}),O(2^{\gamma^{2}n}/(\gamma^{4}n))]$, while Theorem 3 holds for any $\gamma$ and any $s\in[\Omega(1/\gamma^{2+\varepsilon}),O(2^{n}/n)]$. While incomparable, we note the former region is extremely restrictive, and does not include the common setting of $\gamma=\tfrac{1}{n}$. Meanwhile, the latter region contain almost all possible $(\gamma,s)$ pairs. The latter region is short of subsuming the former interval only by an arbitrarily small polynomial factor of $\gamma^{-\varepsilon}$. Removing this $\varepsilon$-slack remains open.
@@
These two points allow our theorem to remain meaningful even for $\delta=o(1)$. Interestingly, the second point shows that the $s\to s\gamma^{2}$ decay is unavoidable even if we permit $H$ to be of density $o(\delta)$.
In summary, our Theorem 3 is stronger in the regime $s\geq 1/\gamma^{2+\varepsilon}$ or $\gamma=O(1/\sqrt{n})$, but gives inferior bounds when $1/\gamma^{2}\leq s\leq(1/\gamma)^{2+\varepsilon}$ and $\gamma=\Omega(1/\sqrt{n})$. [^4] For general $s=\Omega(1/\gamma^{2})$, we get approximating circuits of size $O\left(s\gamma^{2}\cdot\frac{\log s}{\log(s\gamma^{2})}\right)$. See Remark 1. We also note that 1 remains open. In particular, it would be interesting to pin down the optimal dependence of the circuit-size decay on $\delta$.
Our main technical lemma is the strengthening of a beautiful result of Andreev, Clementi, and Rolim [1], which shows that arbitrary Boolean functions can be approximated by small-size circuits.
@@
This result was proven by Andreev, Clementi, and Rolim [1] in the case where $H$ is uniform over $\{0,1\}^{n}$. A short probabilistic argument proves that the circuit size cannot be improved [1, Theorem 4.1].
@@ 1.1 Proof Overview
Although the weaker version of Theorem 4 suffices, it is our impression that the result of [1] is not as well known to the community as it should be. It is a complete resolution to a very natural question in circuit complexity (it is the “approximate version” of Lupanov’s theorem [13]), and the ideas behind the construction are quite useful (e.g., they appear to have been rediscovered in the construction of covering codes of Rabani and Shpilka [15]). This is potentially due to the paper being quite technical and terse, as well as evading search engines. For this reason, we hope to bring attention to this result by giving an exposition and simpler, self-contained proof of Theorem 4 in the appendix, assuming a basic consequence of the fourth-moment method [3] and the existence of asymptotically good codes encodable by linear-size circuits [17]. We now provide a proof overview below.
@@ 1.1.1 Overview of Theorem 4
For ease of exposition, we give an overview of Theorem 4 when $H$ is fixed to be uniform over $\{0,1\}^{n}$. Extending the given arguments to arbitrary distributions $H$ just requires a couple of extra modifications. This overview is morally the same as one provided by Trevisan [20], but perhaps with a slightly different point of view in the latter half.
@@
Unfortunately, a fully random truth table can only be generated by maximally sized circuits. However, we could hope to use a pseudorandom string instead. The only property used about the randomness of $f$ was that its truth table had square-root anticoncentration on subcubes. It turns out $4$-wise uniform strings have square-root anticoncentration with constant probability [3], motivating us to look at this primitive. Implementing the usual 4-wise uniform-generator construction naively in a circuit immediately gives a distribution $\mathcal{C}$ over circuits of size $O(n^{2})$ such that the truth table of $C\sim\mathcal{C}$ is a $4$-wise independent string. With more effort, one can get a distribution over $O(n)$-size circuits, which is optimal. By the fourth-moment method [3], we can argue that the average number of subcubes of $f\oplus\mathcal{C}$ with square-root anticoncentration is at least a constant proportion. Fixing $C\in\mathcal{C}$ that achieves this average, it follows that there is a function $g$ on the first $k$ bits that approximates $f\oplus C$ well by our previous analysis. Consequently, $C\oplus g$ is a circuit of size $O(2^{k}/k+n)$ that approximates $f$ well, as desired.
@@ 2 Preliminaries
Definition 1 ($k$-wise uniformity [8]).
@@
$\Pr_{x\sim X}\left[\left|\sum_{i=1}^{n}v_{i}x_{i}\right|\geq\sqrt{\frac{\sum_{i=1}^{n}v_{i}^{2}}{3}}\right]\geq\frac{2}{11}.$
@@
This is a $4$-wise uniform generator (see [8, Theorem 2.2]). Notice that, as a function of $x$, $G(s)_{x}$ is an evaluation of a degree-$3$ polynomial, which can be done in a circuit of size $O(n^{2})$ by grade-school multiplication (better multiplication algorithms are known, but this suffices). ∎
This theorem suffices to prove Theorem 3. A technical contribution of [1], and a key ingredient behind Theorem 4, is a $4$-wise uniform generator that can be locally computed in linear circuit size, which is the best one could hope for.
Theorem 7 ([1]).
@@ 3 Tightness of Impagliazzo’s Hardcore Lemma
Let $G$ be the $4$-wise uniform generator guaranteed by the hypothesis. For a subcube $c$, denote the indicator variable
$I_{c}(s)\coloneqq\mathbf{1}\left(\left|\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)(-1)^{f(c\circ y)+G(s)_{c\circ y}}\right|\geq\sqrt{\frac{\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)^{2}}{3}}\right).$
By Theorem 5, we have that for each $c$ and random $s$, $\Pr_{s\sim\{0,1\}^{m}}[I_{c}(s)=1]\geq\frac{2}{11}.$ Hence,
@@
and by an averaging argument there exists a choice of $s$ such that $\Pr_{c\sim H^{\prime}}[I_{c}(s)=1]\geq 2/11$. Fix such an $s$. Now for any $c$ with $I_{c}(s)=1$, we can use Cauchy-Schwarz to lower bound
$\left|\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)(-1)^{f(c\circ y)+G(s)_{c\circ y}}\right|\geq\sqrt{\frac{\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)^{2}}{3}}\geq\sqrt{\frac{(1/2^{\ell})\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)}{3}}=\sqrt{\frac{1}{2^{\ell}\cdot 3}}$
Now define $h:\{0,1\}^{n-\ell}\to\{0,1\}$ by the map
@@
$\geq\Pr_{c\sim H^{\prime}}[I_{c}(s)]\cdot\mathbb{E}_{c\sim H^{\prime}}\left[\left|\sum_{y\in\{0,1\}^{\ell}}H_{c}(y)(-1)^{f(c\circ y)+G(s)_{y}}\right|:I_{c}(s)=1\right]$
$\geq\frac{2}{11}\cdot\sqrt{\frac{1}{2^{\ell}\cdot 3}}\geq 2\gamma.$
Hence,
@@
We will first prove the first item. Pick a function $g:\{0,1\}^{k}\to\{0,1\}$ uniformly at random, and define $f(x)\coloneqq g(x_{\leq k})$ to be $g$ on the first $k$ bits of the input. By a Chernoff bound, we know that for a fixed circuit $C:\{0,1\}^{k}\to\{0,1\}$ of size $\leq s$ and $\delta<0.49$,
$\Pr_{f}\left[\Pr_{x}[g(x)=C(x)]\geq 1-\delta\right]\leq 2^{-\Omega(2^{k})}.$
Set $2^{k}=\Omega(s\log s)$. Taking a union bound over all circuits of size $s$ (of which there are $s^{O(s)}$) we note that, with probability at most $s^{O(s)}2^{-\Omega(2^{k})}<1$, $g$ can be $(1-\delta)$-approximated by a size-$s$ circuit. Fix a $g$ that cannot. We now show $f$ cannot be approximated by any circuit $C:\{0,1\}^{n}\to\{0,1\}$ of size $s$. Restricting the last nkn-k bits of $C$ will result in a circuit on the first $k$ bits of size at most $s$. Hence, we can deduce
$\Pr_{x}[f(x)=C(x)]=\mathbb{E}_{x_{>k}}\left[\Pr_{x_{\leq k}}[g(x_{\leq k})=C(x_{\leq k},x_{>k})]\right]<1-\delta$
as desired.
We will now prove this $f$ satisfies the second item. Consider an arbitrary distribution $H$ over $\{0,1\}^{n}$. The marginal distribution of $H$ on the first $k$ bits will be some distribution $H^{\prime}$ over $\{0,1\}^{k}$. As $\gamma>s^{-1/2}>27\cdot 2^{-k/2}$ for large enough $n$, Lemma 1 and Theorem 6 give a circuit $C:\{0,1\}^{k}\to\{0,1\}$ of size $O\left(\frac{\gamma^{2}2^{k}}{\log(\gamma^{2}2^{k})}+k^{2}\right)$ such that $\Pr_{x\sim H^{\prime}}[g(x)=C(x)]\geq\frac{1}{2}+\gamma.$ Letting $C^{\prime}:\{0,1\}^{n}\to\{0,1\}$ be the circuit that applies $C$ to the first $k$ bits, we note that
$\Pr_{x\sim H}[f(x)=C^{\prime}(x)]=\Pr_{x\sim H^{\prime}}[g(x)=C(x)]\geq\frac{1}{2}+\gamma.$
As $C^{\prime}$ has the same size as $C$, and $\frac{\gamma^{2}2^{k}}{\log(\gamma^{2}2^{k})}+k^{2}=O\left(s\gamma^{2}\cdot\left(\frac{\log s}{\log(\gamma^{2}s)}\right)+\log^{2}s\right)=O_{\varepsilon}(s\gamma^{2})$ when $s\geq 1/\gamma^{2+\varepsilon}$, $C^{\prime}$ is a size-$O_{\varepsilon}(s\gamma^{2})$ circuit that $\gamma$-correlates with $f$ over $H$. Consequently, $f$ is a $k$-junta that cannot be $(1-\delta)$-approximated by a size-$s$ circuit, but, over any $H$, $\gamma$-correlates with a size-$O_{\varepsilon}(s\gamma^{2})$ circuit, as desired. ∎